# Three dimensional end-to-end simulation for kilonova emission from a black-hole neutron-star merger

### v1

@@ 5 Results: KN light curves
Interestingly, the gz-band emission observed from b): $135^{\circ}\leq\phi<180^{\circ}$ becomes slightly brighter in the equatorial direction than in the polar direction after 1 d for the present BH-NS model. This brightness dependence on the latitudinal direction is opposite compared to the BNS KN models, for which the emission becomes brighter in the polar direction [^2] However, it should be noted that here we do not consider the impact that the short GRB jet might have on the polar ejecta and on the KN emission (see Hamidani et al. (2024)).. The same latitudinal-angle dependence is also found for the emission observed in the direction of the ejecta bulk motion for the model in Darbha et al. (2021) in which the radioactive heating rate agrees with our model (model H4). The brighter emission in the equatorial plane is explained by the enhancement of the radiation energy flux due to the Doppler effect induced by the bulk motion of the ejecta. It is also important to have the aspect ratio of the dynamical ejecta close to unity to realize the present latitudinal-angle dependence of the emission brightness: otherwise the Doppler effect can be obscured by the suppression of the emission due to the decrease in the projected area toward the observer for the case that the dynamical ejecta have a more oblate shape (see  Darbha et al. (2021) for the discussion).
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 10: Comparison of the isotropically equivalent bolometric luminosities (left) and gzK-band light curves (right) between the fiducial (the same as in Figures 8 and 9) and axisymmetrized (labeled as “2D model”) models. The top and bottom panels denote the light curves observed from $0^{\circ}\leq\theta\leq 28^{\circ}$ and $82^{\circ}\leq\theta\leq 90^{\circ}$, respectively. The solid and dashed curves denote the light curves of the axisymmetrized model and those for the fiducial model observed from b): $135^{\circ}\leq\phi<180^{\circ}$, respectively. The dotted curves in the upper and bottom panels denote the light curves of the fiducial model observed from d): $315^{\circ}\leq\phi<360^{\circ}$ and a): $45^{\circ}\leq\phi<90^{\circ}$, respectively.
To clarify the RT effect of the non-axisymmetric ejecta geometry, we perform a RT simulation for an axisymmetrized ejecta profile. The axisymmetrized ejecta profile is generated by averaging over the rest-mass density, specific internal energy, elemental abundances, and radioactive heating rate profiles obtained by the HD simulation at $t=0.1\,{\rm d}$ with respect to the longitudinal direction. Note that the volume and mass in each grid cell are used as the weights of the average for the rest-mass density and the latter three quantities, respectively. Figure 10 compares the isotropically equivalent bolometric luminosities and gzK-band light curves observed from the polar ($0^{\circ}\leq\theta\leq 28^{\circ}$) and equatorial ($82^{\circ}\leq\theta\leq 90^{\circ}$) directions between the axisymmetrized and fiducial models (the same as in Figures 8 and 9). For the fiducial model, the light curves observed from b): $135^{\circ}\leq\phi<180^{\circ}$, d): $315^{\circ}\leq\phi<360^{\circ}$ (for the polar view), and a): $45^{\circ}\leq\phi<90^{\circ}$ (for the equatorial view) are shown.
@@
The difference in the brightness between the axisymmetrized model and fiducial model is more pronounced for the emission observed from the equatorial direction. For the equatorial view, the bolometric luminosity observed from the longitudinal direction of the brightest region (‘b)’: $135^{\circ}\leq\phi<180^{\circ}$) is brighter approximately by a factor of 1.5–2 for the fiducial model than that for the axisymmetrized model. The gzK-band magnitudes observed from the same direction are also brighter than those for the axisymmetrized model by $\sim 1\,{\rm mag}$. On the other hand, the KN brightness observed from the longitudinal direction of the faintest region (‘a)’: $45^{\circ}\leq\phi<90^{\circ}$) is comparable to or slightly fainter for the fiducial model than that for the axisymmetrized model. This discrepancy in the equatorial brightness between the fiducial and axisymmetrized models is due to the difference in the ejecta aspect ratio: as a consequence of the longitudinal average, the polar projected area of the ejecta is larger than that in the equatorial direction for the axisymmetrized model, which makes photons to preferentially diffuse in the polar direction and thus the equatorial brightness to be fainter.
@@ 5.4 Comparison with various BNS KN models
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Figure 11: Comparison of the bolometric and gzK-band light curves among the present BH-NS KN model and various BNS KN models. For BNS KN models, three cases are shown: a case in which the remnant MNS survives for a short time (the dashed curves; SFHo-125145, Kiuchi et al. (2022); Fujibayashi et al. (2023); Kawaguchi et al. (2023)), a case in which the remnant MNS survives for a long time (the dash-dotted curves; DD2-135135,  Fujibayashi et al. (2020b); Kawaguchi et al. (2022)), and a case in which large-scale magnetic field significantly plays a role in the long-surviving remnant MNS (the dotted curves; MNS75a,  Shibata et al. (2021); Kawaguchi et al. (2022)). Note that the light curve for model SFHo-125145 in the top-right panel is below the plot range.
@@
Figure 11 compares the gzK-band light curves among the present BH-NS KN model and various BNS KN models obtained in our previous studies (Kawaguchi et al., 2021, 2022, 2023). For BNS KN models, three cases are shown as representative: a case in which the remnant massive NS (MNS) survives for a short time (the dashed curves; SFHo-125145, Kiuchi et al. (2022); Fujibayashi et al. (2023); Kawaguchi et al. (2023)), a case in which the remnant MNS survives for a long time (the dash-dot curves; DD2-135135,  Fujibayashi et al. (2020b); Kawaguchi et al. (2022)), and a case in which large-scale magnetic field significantly plays a role in the long-surviving remnant MNS (the dotted curves; MNS75a,  Shibata et al. (2021); Kawaguchi et al. (2022)). We note that the BNS KN models are obtained by imposing axisymmetry in all the post-merger NR simulations, subsequent HD simulations, and RT simulations. For the BH-NS model, we show the light curves observed from b): $135^{\circ}\leq\phi<180^{\circ}$, d): $315^{\circ}\leq\phi<360^{\circ}$ (for the polar view), and a): $45^{\circ}\leq\phi<90^{\circ}$ (for the equatorial view), which represent the longitudinal directions for the brightest and faintest emission, respectively.
@@
The gz-band emission observed from the polar direction ($0^{\circ}\leq\theta\leq 20^{\circ}$) for the present BH-NS KN model is by 0.5–1 mag brighter than that for the BNS models in which the remnant MNS survives only for a short time ($<10\,{\rm ms}$, SFHo-125145), but is by $\approx$ 1 mag fainter than that for the BNS models in which the remnant MNSs survive for a long time ($>1\,{\rm s}$, DD2-135135 and MNS75a). On the other hand, the gz-band emission observed from the equatorial direction ($86^{\circ}\leq\theta\leq 90^{\circ}$) is comparably bright or brighter than those for the BNS models in which remnant MNSs survive for a long time except for the z-band emission of model MNS75a. This is due to the fact that the BNS KN models show stronger latitudinal viewing-angle dependence than the BH-NS KN model and become significantly faint in the equatorial view. The difference in the latitudinal viewing-angle dependence reflects the fact that the dynamical ejecta are the primary source of the emission in the optical wavelength for the BH-NS model, while for the BNS models, the post-merger ejecta are main source of the emission and the dynamical ejecta are mostly acting as the opacity source rather than the emission source (lanthanide curtain effect; Kasen et al. 2015; Kawaguchi et al. 2018, 2020; Bulla 2019; Zhu et al. 2020; Darbha & Kasen 2020; Korobkin et al. 2021). We note that, in the BNS cases, enhancement of the brightness due to the Doppler effect is obscured by the latitudinal-angle dependence of the emission induced by the angle-dependent opacity of the dynamical ejecta.
@@ 6 Summary and Discussions
In this paper, we studied the long-term evolution of the matter ejected in a BH-NS merger by employing the results of the NR simulation and nucleosynthesis calculation, in which both dynamical and post-merger ejecta are followed consistently. In particular, we employed the results for the merger of a $1.35\,M_{\odot}$ NS and $5.4\,M_{\odot}$ BH with the dimensionless spin of 0.75. We confirmed the finding in the previous studies that, thermal pressure induced by radioactive heating in the ejecta could significantly modify the morphology of the ejecta. In our studied case of a BH-NS binary, the dynamical ejecta expand significantly and the aspect ratio becomes close to unity with the fine structure being smeared out in the presence of radioactive heating. On the other hand, the post-merger ejecta were compressed and confined in the region with the radial velocity $\lesssim 0.05\,c$ due to the significant expansion of the dynamical component.
@@
We then computed the KN light curves employing the ejecta profile obtained by the HD simulation of the ejecta matter. We found that our present BH-NS model results in KN light curves that are fainter but longer lasting than those observed in AT2017gfo, reflecting the fact that the emission is primarily powered by the lanthanide-rich massive dynamical ejecta. The optical-band emission is comparable to or fainter than those for the various BNS models obtained in our previous studies. While the peak brightness of the near-infrared emission is also comparable to the BNS models, the time-scale maintaining the brightness is much longer, and the emission comparable to the peak brightness within 1 mag is sustained for more than two weeks for the BH-NS model. The wide-field infrared observations with the ground-based telescopes, such as VISTA (Ackley et al., 2020), WINTER (Frostig et al., 2022) and PRIME (Kondo et al., 2023), can detect such bright infrared KN emission up to $\approx 14$ d if the distance to the event is within 150 Mpc since the K-band emission will be apparently brighter than 21 mag for all the viewing angles. However, the field of views of infrared telescopes are typically not as large as those for the optical telescopes for a given sensitivity (Nissanke et al., 2013). Therefore, a tight constraint of the localization area by the GW data analysis or the follow-up observation within $\approx 1$ d in the optical bands is crucial to detect the KN emission unless the event occurs as close as in the case of AT2017gfo. Once a KN with long-lasting near-infrared emission is found, follow-up observations in the radio band may also be useful to support the presence of massive dynamical ejecta, by finding the synchrotron radio flares with the relatively delayed peak time of $\sim 10\,{\rm yr}$ (Kyutoku et al., 2013).
@@
We found that the non-axisymmetric geometry of the ejecta induces various interesting radiative-transfer effects in the viewing-angle dependence of the KN emission. In particular, we found the Doppler effect induced by the bulk velocity of the ejecta to the emission, which is pointed out by Fernández et al. (2017) and Darbha et al. (2021), is in fact present. Due to this effect, the optical light curves observed from the direction of the bulk ejecta motion show a slightly inverted latitudinal angle dependence to those found in the BNS models: The optical-band emission observed from b): $135^{\circ}\leq\phi<180^{\circ}$ becomes slightly brighter in the equatorial direction than in the polar direction for the present BH-NS model. Since the KNe emission becomes fainter in the equatorial direction than in the polar direction for BNS mergers, our results suggest that, for the edge-on view, the KN emission for BH-NS mergers can be brighter in the optical-band than that of BNS mergers.
@@
Our results indicate that the long-lasting near-infrared emission is the key to distinguish the types of progenitors by the KN observation. If the K-band emission of which brightness comparable to its peak is maintained for more than two weeks, it may indicate that the progenitor is a BH-NS merger with massive ejecta formation. This is consistent with our finding in the previous study (Kawaguchi et al., 2020). On the other hand, only from the optical emission, the BH-NS KN light curves can be similar to those associated with BNS mergers, and hence, it may be difficult to infer the information of the progenitor. We should note that the ejecta mass and hence the brightness of the KN of BH-NS mergers can have large variety depending on the binary parameters, such as the BH and NS masses, BH spin, and NS radius (Rosswog, 2005; Shibata & Taniguchi, 2008; Etienne et al., 2009; Lovelace et al., 2013; Kyutoku et al., 2015; Foucart et al., 2018), as well as on the adopted EOS (Hayashi et al., 2023). We also note that the assumption of LTE employed in our RT simulation will not be valid in the region where the rest-mass density has been significantly dropped. It is implied in Hotokezaka et al. (2021); Pognan et al. (2022) that the matter temperature can be higher than that estimated under the assumption of LTE if non-LTE effects take place. In such a case, the infrared emission can be dimmer with the combination of the suppression of the neutral and poorly ionized atoms (Kawaguchi et al., 2021, 2022, 2023). Hence, we should note that the absence of the long-lasting bright near-infrared emission does not necessarily rule out the possibility that the progenitor of the observed KN is a BH-NS merger.
@@
So far, four candidates have been reported for BH-NS GW events: GW190814 (Abbott et al., 2020), GW200105/GW200115 (Abbott et al., 2021), and GW230529 (The LIGO Scientific Collaboration et al., 2024a). Among them, according to the inferred masses and spins of the binary, the latest GW event, GW230529, was most likely to be accompanied by EM counterparts. Unfortunately, the EM counterpart was not found in GW230529 due to the poorly constrained sky localization, although the luminosity distance to the event was relatively close ($201^{+102}_{-96}\,{\rm Mpc}$ with the error bar being the 90% credible intervals). Nevertheless, the discovery of this system increases the expected rate of the GW detection of a BH-NS merger with EM counterparts in the future. For this event, (under the assumption that this event was a BH-NS merger) the ratio of the BH mass to the NS mass and the dimensionless BH spin were less than $\approx 4$ and larger than $\approx 0$, respectively. For such a case of a BH-NS merger, the ratio of the post-merger ejecta mass to the dynamical one can be larger compared to the BH-NS model which we studied in this paper (Hayashi et al., 2021). This indicates that the resulting KN may become bluer than the present result, while it is not always trivial since the long-term hydrodynamics evolution of ejecta may also differ. Hence, the systematic study on the KNe for various configurations of BH-NS binaries would be crucial to quantitatively interpret the EM observational data in the future.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Figure 12: Comparison between the present BH-NS KN model and GRB KN candidates. The solid and dashed curves denote the polar light curves in the observer frame ($0^{\circ}\leq\theta\leq 28^{\circ}$ ) for the present BH-NS KN model observed from b): $135^{\circ}\leq\phi<180^{\circ}$ and d): $315^{\circ}\leq\phi<360^{\circ}$, respectively. The square and triangle symbols denote, respectively, the observed magnitudes and upper-limits of the optical and near-infrared counterparts of GRBs taken from Jin et al. (2016, 2015); Yang et al. (2015); Berger et al. (2013); Tanvir et al. (2013); Lamb et al. (2019); Troja et al. (2019); Rastinejad et al. (2022); Troja et al. (2022); Gompertz et al. (2023); Levan et al. (2024). The afterglow models which broadly reproduce the models in the literature are also plotted in the dotted curves.
@@
There are a number of KN candidates reported which are associated with the observation of GRBs: GRB050709 (Jin et al., 2016), GRB060614 (Jin et al., 2015; Yang et al., 2015), GRB130603B (Berger et al., 2013; Tanvir et al., 2013), GRB160821B (Lamb et al., 2019; Troja et al., 2019), GRB211211A (Rastinejad et al., 2022; Troja et al., 2022; Gompertz et al., 2023), and GRB230307A (Levan et al., 2024). In Figure 12, we compare our present BH-NS KN model with these observational data. The optical and near-infrared brightness of KN candidates found in GRB211211A and GRB230311 are comparable to that of AT2017gfo. We find that our present BH-NS KN model is too faint to explain the optical brightness of these KN candidates at a few days, while the K-band emission of our present BH-NS KN model after 4 d is too bright to be consistent with the later time upper limits. Our present BH-NS KN model is also too faint in the optical bands to explain the KN candidates found in GRB050709, GRB060614, and GRB160821B after $\approx 2$ d. The K-band brightness of GRB160821B at 4.3 d is comparable to that of our present BH-NS KN model, while the BNS model in which the remnant MNS survives for a long time ($>1\,{\rm s}$, DD2-135135 in Figure 11) also has comparable K-band brightness at that epoch. Interestingly, despite the bright and long-lasting K-band emission, our present BH-NS model has fainter H-band emission at $\sim 10\,{\rm d}$ than that observed in GRB130603B. This is due to the fact that the KN of our present BH-NS model is very red and the peak of the red-shifted spectrum is located in the wavelength longer than the H-band at that epoch. In summary, currently we do not find a KN candidate that can be only explained by our present BH-NS model with significant dynamical ejecta formation. However, as we mentioned above, we cannot rule out the possibility that some of those KN candidates are KNe associated with BH-NS mergers, since BH-NS KNe can have a large diversity reflecting the variety of binary parameters as well as the adopted EOS.

### v2

@@ 5 Results: KN light curves
Interestingly, the gz-band emission observed from b): $135^{\circ}\leq\phi<180^{\circ}$ becomes slightly brighter in the equatorial direction than in the polar direction after 1 d for the present BH-NS model. This brightness dependence on the latitudinal direction is opposite compared to the BNS KN models, for which the emission becomes brighter in the polar direction. The same latitudinal-angle dependence is also found for the emission observed in the direction of the ejecta bulk motion for the model in Darbha et al. (2021) in which the radioactive heating rate broadly agrees with our model (model H4). The brighter emission in the equatorial plane is explained by the enhancement of the radiation energy flux due to the Doppler effect induced by the bulk motion of the ejecta. It is also important to have the aspect ratio of the dynamical ejecta close to unity to realize the present latitudinal-angle dependence of the emission brightness: otherwise the Doppler effect can be obscured by the suppression of the emission due to the decrease in the projected area toward the observer for the case that the dynamical ejecta have a more oblate shape (see  Darbha et al. 2021 for the discussion).
@@
Figure 11: Comparison of the isotropically equivalent bolometric luminosities (left) and gzK-band light curves (right) between the fiducial (the same as in Figures 9 and 10) and axisymmetrized (labeled as “2D model”) models. The top and bottom panels denote the light curves observed from $0^{\circ}\leq\theta\leq 28^{\circ}$ and $82^{\circ}\leq\theta\leq 90^{\circ}$, respectively. The solid and dashed curves denote the light curves of the axisymmetrized model and those for the fiducial model observed from b): $135^{\circ}\leq\phi<180^{\circ}$, respectively. The dotted curves in the upper and bottom panels denote the light curves of the fiducial model observed from d): $315^{\circ}\leq\phi<360^{\circ}$ and a): $45^{\circ}\leq\phi<90^{\circ}$, respectively.
To clarify the RT effect of the non-axisymmetric ejecta geometry, we perform a RT simulation for an axisymmetrized ejecta profile. The axisymmetrized ejecta profile is generated by averaging over the rest-mass density, specific internal energy, elemental abundances, and radioactive heating rate profiles obtained by the HD simulation at $t=0.1\,{\rm d}$ with respect to the longitudinal direction. Note that the volume and mass in each grid cell are used as the weights of the average for the rest-mass density and the latter three quantities, respectively. Figure 11 compares the isotropically equivalent bolometric luminosities and gzK-band light curves observed from the polar ($0^{\circ}\leq\theta\leq 28^{\circ}$) and equatorial ($82^{\circ}\leq\theta\leq 90^{\circ}$) directions between the axisymmetrized and fiducial models (the same as in Figures 9 and 10). For the fiducial model, the light curves observed from b): $135^{\circ}\leq\phi<180^{\circ}$, d): $315^{\circ}\leq\phi<360^{\circ}$ (for the polar view), and a): $45^{\circ}\leq\phi<90^{\circ}$ (for the equatorial view) are shown.
@@
The difference in the brightness between the axisymmetrized model and fiducial model is more pronounced for the emission observed from the equatorial direction. For the equatorial view, the bolometric luminosity observed from the longitudinal direction of the brightest region (‘b)’: $135^{\circ}\leq\phi<180^{\circ}$) is brighter approximately by a factor of 1.5–2 for the fiducial model than that for the axisymmetrized model. The gz-band magnitudes observed from the same direction are also brighter than those for the axisymmetrized model by $\sim 1\,{\rm mag}$. On the other hand, the KN brightness observed from the longitudinal direction of the faintest region (‘a)’: $45^{\circ}\leq\phi<90^{\circ}$) is comparable to or slightly fainter for the fiducial model than that for the axisymmetrized model. This discrepancy in the equatorial brightness between the fiducial and axisymmetrized models is due to the difference in the ejecta aspect ratio: as a consequence of the longitudinal average, the polar projected area of the ejecta is larger than that in the equatorial direction for the axisymmetrized model, which makes photons to preferentially diffuse in the polar direction and thus the equatorial brightness to be fainter.
@@ 5.4 Comparison with various BNS KN models
Figure 12: Comparison of the bolometric and gzK-band light curves among the present BH-NS KN model and various BNS KN models. For BNS KN models, three cases are shown: a case in which the remnant MNS survives for a short time (the dashed curves; SFHo-125145, Kiuchi et al. 2022; Fujibayashi et al. 2023; Kawaguchi et al. 2023), a case in which the remnant MNS survives for a long time (the dash-dotted curves; DD2-135135,  Fujibayashi et al. 2020b; Kawaguchi et al. 2022), and a case in which large-scale magnetic field significantly plays a role in the long-surviving remnant MNS (the dotted curves; MNS75a,  Shibata et al. 2021; Kawaguchi et al. 2022). Note that the light curve for model SFHo-125145 in the top-right panel is below the plot range. Note also that due to different angle binning setups, the light curves observed from $0^{\circ}\leq\theta\leq 20^{\circ}$ and $86^{\circ}\leq\theta\leq 90^{\circ}$ are used for BNS KN models to compare with the BH-NS KN light curves observed from $0^{\circ}\leq\theta\leq 28^{\circ}$ and $82^{\circ}\leq\theta\leq 90^{\circ}$, respectively. The data points denote the AB magnitudes of AT2017gfo taken from Villar et al. (2017).
@@
Figure 12 compares the gzK-band light curves among the present BH-NS KN model and various BNS KN models obtained in our previous studies (Kawaguchi et al., 2021, 2022, 2023). For BNS KN models, three cases are shown as representative: a case in which the remnant massive NS (MNS) survives for a short time (the dashed curves; SFHo-125145, Kiuchi et al. 2022; Fujibayashi et al. 2023; Kawaguchi et al. 2023), a case in which the remnant MNS survives for a long time (the dash-dot curves; DD2-135135,  Fujibayashi et al. 2020b; Kawaguchi et al. 2022), and a case in which large-scale magnetic field significantly plays a role in the long-surviving remnant MNS (the dotted curves; MNS75a,  Shibata et al. 2021; Kawaguchi et al. 2022).
For model SFHo-125145, the ejecta mass and average velocity are $0.012\,M_{\odot}$ and $0.17\,c$, respectively. The dynamical and post-merger ejecta have approximately the same mass and exhibit approximately spherical and mildly prolate shapes, respectively. While the $Y_{e}$ value of the post-merger ejecta is $\gtrsim 0.3$ and shows a weak spatial dependence, the dynamical ejecta have a significant $Y_{e}$ dependence on the latitudinal angle with the value being higher and lower than 0.3 for $\theta\lesssim 45^{\circ}$ and $\gtrsim 45^{\circ}$, respectively. For model DD2-135135, the ejecta mass and average velocity are $0.065\,M_{\odot}$ and $0.094\,c$, respectively. The morphology and $Y_{e}$ profiles of the dynamical and post-merger ejecta are similar to those in model SFHo-125145, but the dynamical ejecta mass is $1.5\times 10^{-3}\,M_{\odot}$, and the post-merger ejecta dominate the total ejecta mass. For model MNS75a, the same BNS configuration as in model DD2-135135 is considered, but the post-merger ejecta are significantly accelerated by taking into account the magneto-centrifugal effect. As the consequence, the ejecta mass and average velocity are increased compared to those in model DD2-135135 and are $0.084\,M_{\odot}$ and $0.42\,c$, respectively.
We note that the BNS KN models are obtained by imposing axisymmetry in all the post-merger NR simulations, subsequent HD simulations, and RT simulations. For the BH-NS model, we show the light curves observed from b): $135^{\circ}\leq\phi<180^{\circ}$, d): $315^{\circ}\leq\phi<360^{\circ}$ (for the polar view), and a): $45^{\circ}\leq\phi<90^{\circ}$ (for the equatorial view), which represent the longitudinal directions for the brightest and faintest emission, respectively.
@@
The gz-band emission observed from the polar direction ($0^{\circ}\leq\theta\leq 28^{\circ}$) for the present BH-NS KN model is by 0.5–1 mag brighter than that for the BNS models in which the remnant MNS survives only for a short time ($<10\,{\rm ms}$, SFHo-125145), but is by more than $\approx$ 1 mag fainter than that for the BNS models in which the remnant MNS survives for a long time ($>1\,{\rm s}$, DD2-135135 and MNS75a). On the other hand, the gz-band emission observed from the equatorial direction ($82^{\circ}\leq\theta\leq 90^{\circ}$) is comparably bright or brighter than those for the BNS models in which remnant MNSs survive for a long time except for the z-band emission of model MNS75a. This is due to the fact that the BNS KN models show stronger latitudinal viewing-angle dependence than the BH-NS KN model and become significantly faint in the equatorial view. The difference in the latitudinal viewing-angle dependence reflects the fact that the dynamical ejecta are the primary source of the emission in the optical wavelength for the BH-NS model, while for the BNS models, the post-merger ejecta are main source of the emission and the dynamical ejecta are mostly acting as the opacity source rather than the emission source (lanthanide curtain effect; Kasen et al. 2015; Kawaguchi et al. 2018; Kawaguchi et al. 2020; Bulla 2019; Zhu et al. 2020; Darbha & Kasen 2020; Korobkin et al. 2021). We note that, in the BNS cases, enhancement of the brightness due to the Doppler effect is obscured by the latitudinal-angle dependence of the emission induced by the angle-dependent opacity of the dynamical ejecta.
@@ 6 Summary and Discussions
In this paper, we studied the long-term evolution of the matter ejected in a BH-NS merger by employing the results of the NR simulation and nucleosynthesis calculation, in which both dynamical and post-merger ejecta are followed consistently. In particular, we employed the results for the merger of a $1.35\,M_{\odot}$ NS and $5.4\,M_{\odot}$ BH with the dimensionless spin of 0.75. We confirmed the finding in the previous studies that, thermal pressure induced by radioactive heating in the ejecta could significantly modify the morphology of the ejecta. In our studied case for the BH-NS binary, the dynamical ejecta expand significantly and the aspect ratio becomes close to unity with the fine structure being smeared out in the presence of radioactive heating. On the other hand, the post-merger ejecta were compressed and confined in the region with the radial velocity $\lesssim 0.05\,c$ due to the significant expansion of the dynamical component.
@@
We then computed the KN light curves employing the ejecta profile obtained by the HD simulation of the ejecta matter. We found that our present BH-NS model results in KN light curves that are fainter but longer lasting than those observed in AT2017gfo, reflecting the fact that the emission is primarily powered by the lanthanide-rich massive dynamical ejecta. The optical-band emission is comparable to or fainter than those for the various BNS models obtained in our previous studies. While the peak brightness of the near-infrared emission is also comparable to the BNS models, the time-scale maintaining the brightness is much longer, and the emission comparable to the peak brightness within 1 mag is sustained for more than two weeks for the BH-NS model. The wide-field infrared observations with the ground-based telescopes, such as VISTA (Ackley et al., 2020), WINTER (Frostig et al., 2022) and PRIME (Kondo et al., 2023), can detect such bright infrared KN emission up to $\approx 14$ d if the distance to the event is within 150 Mpc since the K-band emission will be apparently brighter than 21 mag for all the viewing angles. However, the field of view of infrared telescopes is typically not as large as those for the optical telescopes for a given sensitivity (Nissanke et al., 2013). Therefore, a tight constraint of the localization area by the GW data analysis or the follow-up observation within $\approx 1$ d in the optical bands is crucial to detect the KN emission unless the event occurs as close as in the case of AT2017gfo.
@@
We found that the non-axisymmetric geometry of the ejecta induces various interesting radiative-transfer effects in the viewing-angle dependence of the KN emission. In particular, we found the Doppler effect induced by the bulk velocity of the ejecta to the emission, which is pointed out by Fernández et al. (2017) and Darbha et al. (2021), is in fact appreciable. Due to this effect, the optical light curves observed from the direction of the bulk ejecta motion show a slightly inverted latitudinal angle dependence to those found in the BNS models: The optical-band emission observed from b): $135^{\circ}\leq\phi<180^{\circ}$ becomes slightly brighter in the equatorial direction than in the polar direction for the present BH-NS model. Since the KNe emission becomes fainter in the equatorial direction than in the polar direction for BNS mergers, our results suggest that, for the edge-on view, the KN emission for BH-NS mergers can be brighter in the optical-band than that of BNS mergers.
@@
Our results indicate that the long-lasting near-infrared emission is the key to distinguish the types of progenitors by the KN observation. If the K-band emission of which brightness comparable to its peak is maintained for more than two weeks, it may indicate that the progenitor is a BH-NS merger with massive ejecta formation. This is consistent with our finding in the previous study (Kawaguchi et al., 2020). On the other hand, the BH-NS KN light curves in the optical band is similar to those associated with BNS mergers, and hence, it may be difficult to infer the information of the progenitor only from the optical emission. We should note that the ejecta mass and hence the brightness of the KN of BH-NS mergers can have large variety depending on the binary parameters, such as the BH and NS masses, BH spin, and NS radius (Rosswog, 2005; Shibata & Taniguchi, 2008; Etienne et al., 2009; Lovelace et al., 2013; Kyutoku et al., 2015; Foucart et al., 2018), as well as on the adopted EOS (Hayashi et al., 2023). We also note that the assumption of LTE employed in our RT simulation will not be valid in the region where the rest-mass density has been significantly dropped. It is implied in Hotokezaka et al. (2021); Pognan et al. (2022) that the matter temperature can be higher than that estimated under the assumption of LTE if non-LTE effects take place. In such a case, the infrared emission can be dimmer with the combination of the suppression of the neutral and poorly ionized atoms (Kawaguchi et al., 2021, 2022, 2023). We should keep in mind that the absence of the long-lasting bright near-infrared emission does not necessarily rule out the possibility that the progenitor of the observed KN is a BH-NS merger.
@@
So far, four candidates have been reported for BH-NS GW events: GW190814 (Abbott et al., 2020), GW200105/GW200115 (Abbott et al., 2021), and GW230529 (The LIGO Scientific Collaboration et al., 2024). Among them, according to the inferred masses and spins of the binary, the latest GW event, GW230529, was most likely to be accompanied by EM counterparts. Unfortunately, the EM counterpart was not found in GW230529 due to the poorly constrained sky localization, although the luminosity distance to the event was relatively close ($201^{+102}_{-96}\,{\rm Mpc}$ with the error bar being the 90% credible intervals). Nevertheless, the discovery of this system increases the expected rate of the GW detection of a BH-NS merger with EM counterparts in the future. For this event, (under the assumption that this event was a BH-NS merger) the ratio of the BH mass to the NS mass and the dimensionless BH spin were less than $\approx 4$ and larger than $\approx 0$, respectively. For such a case of a BH-NS merger, the ratio of the post-merger ejecta mass to the dynamical one can be larger compared to the BH-NS model which we studied in this paper (Hayashi et al., 2021). This indicates that the resulting KN may become bluer than the present result, while it is not always trivial since the long-term hydrodynamics evolution of ejecta may also differ. Hence, the systematic study on the KNe for various configurations of BH-NS binaries would be crucial to quantitatively interpret the EM observational data in the future.
@@
Figure 13: Comparison between the present BH-NS KN model and GRB KN candidates. The solid and dashed curves denote the polar light curves in the observer frame ($0^{\circ}\leq\theta\leq 28^{\circ}$ ) for the present BH-NS KN model observed from b): $135^{\circ}\leq\phi<180^{\circ}$ and d): $315^{\circ}\leq\phi<360^{\circ}$, respectively. The square and triangle symbols denote, respectively, the observed magnitudes and upper-limits of the optical and near-infrared counterparts of GRBs taken from Jin et al. (2016); Jin et al. (2015); Yang et al. (2015); Berger et al. (2013); Tanvir et al. (2013); Lamb et al. (2019); Troja et al. (2019); Rastinejad et al. (2022); Troja et al. (2022); Gompertz et al. (2023); Levan et al. (2024). The afterglow models which broadly reproduce the models in the literature are also plotted in the dotted curves.
@@
There are a number of KN candidates reported which are associated with the observation of GRBs: GRB050709 (Jin et al., 2016), GRB060614 (Jin et al., 2015; Yang et al., 2015), GRB130603B (Berger et al., 2013; Tanvir et al., 2013), GRB160821B (Lamb et al., 2019; Troja et al., 2019), GRB211211A (Rastinejad et al., 2022; Troja et al., 2022; Gompertz et al., 2023), and GRB230307A (Levan et al., 2024). In Figure 13, we compare our present BH-NS KN model with these observational data. The optical and near-infrared brightness of KN candidates found in GRB211211A and GRB230311 are comparable to that of AT2017gfo. We find that our present BH-NS KN model is too faint to explain the optical brightness of these KN candidates at a few days, while the K-band emission of our present BH-NS KN model after 4 d is too bright to be consistent with the later time observations and upper limits. Our present BH-NS KN model is also too faint in the optical bands to explain the KN candidates found in GRB050709, GRB060614, and GRB160821B after $\approx 2$ d. The K-band brightness of GRB160821B at 4.3 d is comparable to that of our present BH-NS KN model, while the BNS model in which the remnant MNS survives for a long time ($>1\,{\rm s}$, DD2-135135 in Figure 12) also has comparable K-band brightness at that epoch. Interestingly, despite the bright and long-lasting K-band emission, our present BH-NS model has fainter H-band emission at $\sim 10\,{\rm d}$ than that observed in GRB130603B. This is due to the fact that the KN of our present BH-NS model is very red and the peak of the red-shifted spectrum is located in the wavelength longer than the H-band at that epoch. In summary, currently we do not find a KN candidate that can be only explained by our present BH-NS model with significant dynamical ejecta formation. However, as we mentioned above, we cannot rule out the possibility that some of those KN candidates are KNe associated with BH-NS mergers, since BH-NS KNe can have a large diversity reflecting the variety of binary parameters as well as the adopted EOS.